8.)A wire carries a steady current of 0.1 Ampere over a period of 20 seconds. What total charge passes through the wire in this time interval? (3 points) 9.)A 12- C charge is located in a 350-N/C electric field. What is the force on the charge? (3points) 10.) Find the equivalent resistance of the circuit shown below. (6 points) 1o0n M- 4,00 Ω 8.00 V
Answer #1

8.

I=rac{Q}{t}

Rightarrow Q=I imes t=0.1 imes 20=2 C

9.

F = QE 12 × 10-6 × 350= 4.2 × 10-3 = 4.2 mN

10.

10.0 Ω 4.00 Ω 1 5,00 Ω 2.00 Ω 3.00 Ω 8.00 V

The effective resistance at 1 is

3 R 0 5 10

10

The effective resistance at 2 is

1 31 R2R1+4 3 10+12 366

R2 = 31

And the total resistance is

R=2+R2 =- + 2 = 4.1352 31